We have detected an issue on the switch for this rack.
Our datacenter and network teams are working on it to restore the service as quickly as possible.
We will keep you updated as soon as we have more information.
08/06/2019 1550Z (1750LT)
The dedicated team is still working on the issue, we will update this status as soon as we know more.
08/06/2019 1930Z (2130LT)
The issue has been solved, all server are now up and running.